If the length of the diagonal of a square is increasing at the rate of 0.2 m / sec, what is the
rate of increase of its area when the side is
2
30 cm?
a) 3 cm2
/ sec b) 6 2 cm2
/ sec c) 3 2 cm2
/ sec d) 6 cm2
/ sec

Grade:

2 Answers

Arun
3 years ago

let the side of square be X cm at any instant of time and p be it's perimeter 

 

Now :- dx/dt = 0.2cm/s 

p=4x

so 

dp/dt = 4*dx/dt

 

dp/dt= 4*0.2=0.8cm/s

Vikas TU
3 years ago
Dear student 
Question is not clear 
Please attach an image, 
We will happy to  help you 
Good Luck
Cheers